OpenStudy (anonymous):

Find the equation of the line that passes through (-3, 2) with a slope of (-4/5). Write your answer in Slope-Intercept Form

OpenStudy (anonymous):

\[y-2=\frac{-4}{5}(x+3)\]

OpenStudy (anonymous):

In slope intercept for we get:\[y=\frac{-4}{5}x-\frac{2}{5}\]

OpenStudy (precal):

y=mx+b 2=(-4/5)(-3)+b 2=(12/5)+b -2/5=b y=(-4/5)x-(2/5)

OpenStudy (muhammad_nauman_umair):

y-2=(-4/5)(x+3 )